# Linearizability of flows by embeddings

### v6

@@ Abstract.
We consider the problem of determining the class of continuous-time dynamical systems that can be globally linearized in the sense of admitting an embedding into a linear system on a finite-dimensional Euclidean space. We solve this problem for dynamical systems having either a compact state space or at least one compact attractor, obtaining necessary and sufficient conditions for the existence of linearizing $C^{k}$ embeddings for $k\in\mathbb{N}_{\geq 0}\cup\{\infty\}$. Corollaries include (i) several checkable necessary conditions for global linearizability and (ii) extensions of the Hartman-Grobman and Floquet normal form theorems beyond the classical settings. Overall, our results reveal relationships between linearizability, symmetry, topology, and invariant manifold theory. In particular, these relationships yield fundamental limitations and capabilities of algorithms from the “applied Koopman operator theory” literature.
@@ 1.1. Context
for all $t\in\mathbb{R}$. In this paper we study the existence of linearizing maps that are also $C^{k}$ embeddings, where $k\in\mathbb{N}_{\geq 0}\cup\{\infty\}$. [^1] Standard definitions: a map is $C^{0}$ if it is continuous, and is $C^{k}$ with $k\in\mathbb{N}_{\geq 1}\cup\{\infty\}$ if it is $C^{0}$ and has $C^{0}$ partial derivatives of all orders less than $k+1$ when expressed in local coordinates. A $C^{0}$ embedding (or topological embedding) is a homeomorphism onto its image in the subspace topology [Lee11, p. 54]. An immersion is a $C^{1}$ map with an injective derivative at each point. A $C^{k}$ embedding with $k\in\mathbb{N}_{\geq 1}\cup\{\infty\}$ is a $C^{k}$ map that is both an immersion and a $C^{0}$ embedding [Hir94, p. 21]. Equivalently, a $C^{k\geq 1}$ map is a $C^{k}$ embedding if its image is a $C^{k}$ embedded submanifold and the map is a diffeomorphism from its domain to its image. Note that $n\geq\dim M$ if $F\colon M\to\mathbb{R}^{n}$ is a $C^{0}$ embedding, i.e., embeddings always increase or preserve dimension.
Poincaré initiated the study of locally-defined, dimension-preserving linearizing embeddings (homeomorphisms and diffeomorphisms) near an equilibrium point [Poi79, Poi90], leading to classical linearization theorems including those of Poincaré-Siegel [Poi79, Sie42, Sie52], Sternberg [Ste57, Ste58], Hartman-Grobman [Har60a, Gro59] and Hartman [Har60b], as well as recent results like those of Newhouse [New17] and Kvalheim and Revzen [KR21]. Extensions to locally-defined, dimension-preserving, partial linearizations (in transverse directions) near general invariant manifolds were given in theorems including those of Pugh and Shub [PS70], Robinson [Rob71], Takens [Tak71], Sell [Sel83, Sel84], and Sakamoto [Sak94]. Such local linearization results were extended to global ones valid on the entire basin of attraction of a stable hyperbolic equilibrium or limit cycle by Lan and Mezić [LM13] (see also [KR21]) and, more generally, on the entire basin of attraction of a stable normally hyperbolic invariant manifold by Eldering, Kvalheim, and Revzen [EKR18].
To our knowledge, Carleman was the first to study linearizing embeddings without the dimension-preserving assumption [Car32, KS91]. Conceptually, such linearizing embeddings identify nonlinear systems with invariant subsets of linear systems. Currently, such embeddings are viewed as being of paramount importance in the rapidly developing field of applied Koopman operator theory [BBKK22, Mez24]. In fact, numerical approximations of linearizing embeddings are currently being sought via machine learning and “big data”-driven tools ([BBKK22, §1.3, §5], [Mez24]) like deep neural networks, as in Lusch, Brunton, and Kutz [LKB18]. Linearizing embeddings have also been studied in the control theory literature (see below) [CFI83, LM86, Bel22a, Bel22b, BC23, LOS23a, LOS23b, KB24], and they are somewhat related to “polynomial flows” [vdE94, p. 671]. However, linearizing embeddings do not generally exist, raising the fundamental
@@
We answer this question for nonlinear systems having either a compact state space or at least one compact attractor, by obtaining necessary and sufficient conditions.
We emphasize that linearizing $C^{k}$ embeddings are conceptually natural objects of study, since they are precisely the maps $C^{k}$-identifying nonlinear systems with invariant subsets of linear systems. Stated differently, they are precisely the $C^{k}$ linearizing maps having $C^{k}$ left inverses defined on neighborhoods of their images. However, previous literature has considered different classes of maps, as follows.
Liu, Ozay, and Sontag [LOS23a, LOS23b] obtained a necessary condition for linearizability by injective $C^{0}$ maps, which are somewhat more general than $C^{0}$ embeddings, but equivalent in various situations including those considered in the present paper. On the other hand, other control theory literature has focused on sufficient and/or necessary conditions for the existence of $C^{k\geq 1}$ embeddings $F\colon M\to\mathbb{R}^{n}$ satisfying the extra requirements that $M=\mathbb{R}^{m}$ and $F\colon\mathbb{R}^{m}\to\mathbb{R}^{n}$ has a linear left inverse $\mathbb{R}^{n}\to\mathbb{R}^{m}$, while only satisfying a weaker “linearizing” property than we require. Namely, the linear projection is merely required to send linear trajectories starting in the image of $F$ to nonlinear trajectories in $M$, with the image of $F$ not necessarily invariant under the linear flow. Belabbas and Chen called such embeddings $F$ “super-linearizations” [Bel22a, Bel22b, BC23]. We note that Claude, Fliess, and Isidori showed that a $C^{\infty}$ flow $\Phi$ on $\mathbb{R}^{m}$ is super-linearizable if and only if $\text{span}\{x\mapsto\frac{d^{j}}{dt^{j}}\Phi^{t}(x)|_{t=0}\colon j\in\mathbb{N}\}$ is a finite-dimensional linear subspace of the real vector space of $C^{\infty}$ maps $\mathbb{R}^{m}\to\mathbb{R}^{m}$ [CFI83] (see also Levine and Marino [LM86, Lem. 1]). Ko and Belabbas defined “strong super-linearizations” to be super-linearizations that are also linearizing in our sense [KB24]. Existence of a $C^{k}$ strong super-linearization for a flow $\Phi$ on $\mathbb{R}^{m}$ is equivalent to its group of Koopman operators having a finite-dimensional invariant subspace of $C^{k}$ functions containing the state coordinate projections $(x_{1},\ldots,x_{m})\mapsto x_{i}$, a situation of interest in the applied Koopman operator literature since the work of Brunton et al [BBPK16].
In particular, our necessary and sufficient conditions for the existence of a $C^{k\geq 0}$ linearizing embedding yield necessary conditions for the existence of one of the types of maps described in the preceding paragraph.
@@
Returning to the context of linearizing embeddings addressed in the present paper, we note that our existence question fits into the broader context of “universality” introduced by Tao to study the global existence problem for the Euler and Navier-Stokes equations [Tao17, Tao18, Tao20]. Slightly generalizing Tao’s definition [Tao17, p. 220], we say that a class $\mathcal{C}$ of flows is $C^{k}$-universal with respect to another such class $\mathcal{D}$ if any flow in $\mathcal{D}$ admits an (equivariant) $C^{k}$ embedding into some flow in $\mathcal{C}$ (realizing the former as an invariant subset of the latter). If we fix
[…]
@@
our existence question can be restated equivalently as follows: what is the largest class of flows $\mathcal{D}$ with respect to which $\mathcal{C}$ is $C^{k}$-universal? We answer this question under the additional constraint that $\mathcal{D}\subset\mathcal{E}$, where $\mathcal{E}$ consists of flows having either a compact state space or at least one compact attractor.
@@ 1.2. Illustration of results
One obstruction to the existence of a linearizing $C^{0}$ embedding, if $M$ is connected, occurs if (1) has a compact attractor whose basin is not all of $M$ (see Corollary 6). Hence we are led to study the existence of linearizing embeddings defined on attractor basins or other subsets, i.e., linearizing embeddings of $(X,\Phi)$ where $X\subset M$ is a subset invariant under $\Phi$ (by abuse of notation we still write $\Phi$ for $\Phi|_{\mathbb{R}\times X}$). Such a linearizing embedding $F\colon X\to\mathbb{R}^{n}$ satisfies (2) but is defined on $X$ rather than $M$.
Our results are divided among four cases: linearization by a $C^{k\geq 1}$ embedding when $X$ is an attractor basin or a compact invariant manifold, and linearization by a $C^{0}$ embedding when $X$ is an attractor basin or a compact invariant set. Given the mentioned obstruction (Corollary 6), these cases cover linearizability by $C^{k\geq 0}$ embeddings when there is a compact state space or at least one compact attractor.
@@
Our second main result (Theorem 2 in §2.2) concerns the case that $X$ is a compact invariant set (not assumed to be a manifold). It asserts that $(X,\Phi)$ is linearizable by a $C^{0}$ embedding if and only if $\Phi$ is a $1$-parameter subgroup of a $C^{0}$ torus action that is not too pathological (has finitely many “orbit types”). For example, all flows shown in Figure 1 are linearizable by $C^{0}$ embeddings (Figure 1).
Our third main result (Theorem 3) concerns the case that $X$ is the basin of attraction of an asymptotically stable compact invariant set $A\subset M$ [^2] Actually, $A$ need only be globally asymptotically stable within an invariant set $X\subset M$, as long as $X$ is “locally closed” in $M$ (see Remark 7). It asserts that $(X,\Phi)$ is linearizable by a $C^{0}$ embedding if and only if $(A,\Phi)$ is linearizable by a $C^{0}$ embedding (cf. Theorem 2) and $A$ has $C^{0}$ “asymptotic phase” for $\Phi$. Asymptotic phase as defined in §2.3 generalizes that associated with normally hyperbolic invariant manifolds [Fen74, HPS77, BK94, EKR18], which in turn generalizes that associated with hyperbolic periodic orbits [Hal80, Guc75, Win01]. Theorem 3 further asserts that linearizing $C^{0}$ embeddings (or, more generally, injective $C^{0}$ maps) are always proper and hence “blow up” near the basin boundary $\partial X$, leading to the aforementioned statement (Corollary 6) that $(M,\Phi)$ is not linearizable by a $C^{0}$ embedding if $M$ is connected and there is a non-global compact attractor. Other corollaries of Theorem 3 are extensions of the Hartman-Grobman and Floquet normal form theorems, at the cost of allowing “extra” linearizing coordinates (Corollaries 4, 5).
Our fourth and final main result (Theorem 4) again concerns the case that $X$ is the basin of attraction of an asymptotically stable compact invariant set $A\subset M$, but instead concern linearizability by a $C^{k\geq 1}$ embedding. It asserts the following. If $(X,\Phi)$ is linearizable by a $C^{k}$ embedding, then $A\subset X$ must in fact be a $C^{k}$ embedded submanifold, $(A,\Phi)$ must be linearizable by a $C^{k}$ embedding (cf. Theorem 1), and $A$ must have $C^{k}$ asymptotic phase. Conversely, $(X,\Phi)$ is linearizable by a $C^{k}$ embedding if $A,X,\Phi$ have the preceding properties, $\Phi$ is the flow of a $C^{k}$ vector field (cf. Remark 9), and $\Phi$ is “transversely linearizable” near $A$ in a sense related to normal hyperbolicity (cf. Remarks 11, 12).
The remainder of the paper is organized as follows. Our results and examples are in §2. Proofs are in §3. Closing remarks are in §4.
@@ 2. Results and examples
[…]
@@ 2.1. The smooth compact case
With the aid of Theorem 1, the class of linearizable systems $(X,\Phi)$ as in its statement can be seen to be a subclass of the class of “non-Hamiltonian integrable systems” from the literature [Zun06, Def. 2.6].
@@
Suppose $(X,\Phi)$ is either (i) a single equilibrium point, (ii) a single periodic orbit viewed as a constant-speed rotation on the circle $X=S^{1}$, or (iii) a single quasiperiodic torus viewed as a product of constant-speed rotations on the ($n$-)torus $X=T^{n}=S^{1}\times\cdots\times S^{1}$, a product of circles. It is well-known that all of these cases are linearizable by $C^{\infty}$ embeddings. Let us compare this fact with Theorem 1. In all cases, $(X,\Phi)$ is a $1$-parameter subgroup of a $C^{\infty}$ torus action on $X$: the actions $\Theta$ can be respectively taken to be (i) the trivial action of $T^{1}$ on the equilibrium, (ii) the action of $S^{1}=T^{1}$ on itself, or (iii) the action of $T^{n}$ on itself.
@@
Let $X=S^{2}\subset\mathbb{C}\times\mathbb{R}\approx\mathbb{R}^{3}$ be the unit $2$-sphere centered at the origin. Let $\Theta$ be the $C^{\infty}$ action of the circle $S^{1}\approx\mathbb{R}/\mathbb{Z}$ on $\mathbb{C}\times\mathbb{R}$ defined by $\Theta^{h}(z,s)\coloneqq(e^{2\pi ih}z,s)$. This action leaves $S^{2}$ invariant because $\Theta$ simply rotates the sphere’s latitudinal circles. Thus, Theorem 1 implies that $(X,\Phi)$ with $\Phi^{t}\coloneqq\Theta^{t\mod 1}$ is linearizable by a $C^{\infty}$ embedding. This conclusion can also be seen directly since $\Phi$ is the restriction to $X\subset\mathbb{C}\times\mathbb{R}$ of the linear flow $(t,z,y)\mapsto(e^{2\pi it}z,y)$. In anticipation of Proposition 1, note that $(X,\Phi)$ has two equilibria, each with Hopf index equal to $1$.
@@
Proposition 1 and the Poincaré-Hopf theorem [Mil97, p. 35] imply the following corollary, which is consistent with Examples 2, 3, 4, 5.
@@
The standard Poincaré-Hopf theorem and other results for $C^{\infty}$ manifolds also apply to $C^{1}$ manifolds since any $C^{1}$ manifold has a compatible $C^{\infty}$ structure [Hir94, Thm 2.2.9].
@@
The following corollary follows directly from Corollary 2, Examples 2, 3, 4, 5, and the classification of surfaces [Hir94, Thm 9.3.5].
@@
Theorem 1 gives a necessary and sufficient condition for linearizability by a $C^{k\geq 1}$ embedding. On the other hand, Proposition 1 and Corollaries 2, 3 give necessary conditions for the same, while the following proposition gives a sufficient condition. Proposition 2 is a “rigidity” statement about Koopman eigenfunctions, and its proof is independent of our other results. (Assuming the existence of $F$ below is equivalent to assuming the existence of $n$ nowhere-zero $C^{1}$ Koopman eigenfunctions whose eigenvalues are rationally independent [BBKK22, Sec. 2.2].)

### v7

@@ Abstract.
We consider the problem of determining the class of continuous-time dynamical systems that can be globally linearized in the sense of admitting an embedding into a linear system on a higher-dimensional Euclidean space. We solve this problem for dynamical systems on connected state spaces that are either compact or contain at least one nonempty compact attractor, obtaining necessary and sufficient conditions for the existence of linearizing $C^{k}$ embeddings for $k\in\mathbb{N}_{\geq 0}\cup\{\infty\}$. Corollaries include (i) several checkable necessary conditions for global linearizability and (ii) extensions of the Hartman-Grobman and Floquet normal form theorems beyond the classical settings. Our results open new perspectives on linearizability by establishing relationships to symmetry, topology, and invariant manifold theory.
@@ 1. Introduction
for all $t\in\mathbb{R}$. In other words, $y(t)\coloneqq F(\Phi^{t}(x))$ solves the linear system of equations
$\dot{y}=By.$
In this paper we study the existence of linearizing maps that are also $C^{k}$ embeddings, where $k\in\mathbb{N}_{\geq 0}\cup\{\infty\}$. [^1] Standard definitions: a map is $C^{0}$ if it is continuous, and is $C^{k}$ with $k\in\mathbb{N}_{\geq 1}\cup\{\infty\}$ if it is $C^{0}$ and has $C^{0}$ partial derivatives of all orders less than $k+1$ when expressed in local coordinates. A $C^{0}$ embedding (or topological embedding) is a homeomorphism onto its image in the subspace topology [1, p. 54]. An immersion is a $C^{1}$ map with an injective derivative at each point. A $C^{k}$ embedding with $k\in\mathbb{N}_{\geq 1}\cup\{\infty\}$ is a $C^{k}$ map that is both an immersion and a $C^{0}$ embedding [2, p. 21]. Equivalently, a $C^{k\geq 1}$ map is a $C^{k}$ embedding if its image is a $C^{k}$ embedded submanifold and the map is a diffeomorphism from its domain to its image. Note that $n\geq\dim M$ if $F\colon M\to\mathbb{R}^{n}$ is a $C^{0}$ embedding, i.e., embeddings always increase or preserve dimension.
Remark 1.
One can define linearizing maps and $C^{k}$ embeddings $F\colon M\to\mathbb{C}^{n}$ in exactly the same way. But since $\mathbb{R}^{n}\subset\mathbb{C}^{n}$ and $\mathbb{C}^{n}$ is isomorphic to $\mathbb{R}^{2n}$ as a real vector space, a linearizing $C^{k}$ embedding $M\to\mathbb{C}^{n_{1}}$ exists for some $n_{1}$ if and only if a linearizing $C^{k}$ embedding $M\to\mathbb{R}^{n_{2}}$ exists for some $n_{2}$. Thus, while we choose to emphasize $\mathbb{R}^{n}$-valued linearizing embeddings, the equivalent $\mathbb{C}^{n}$-valued viewpoint should be kept in mind. In particular, this viewpoint is useful for constructing linearizing embeddings given by concatenating eigenfunctions $M\to\mathbb{C}$ of the group of Koopman operators [3, 4, 5, 6, 7].
Conceptually, linearizing embeddings identify nonlinear systems with invariant subsets of linear systems. We ask the fundamental
@@
We answer this question for the class of nonlinear systems on connected state spaces that are either compact or contain at least one nonempty compact attractor, obtaining necessary and sufficient conditions. (Our main results do not actually require a connected state space, but the preceding sentence is a convenient summary.)
@@
These necessary and sufficient conditions will be illustrated in §1.2 after discussing related work. For now, we place them in the broader context of “universality” introduced by Tao to study the global existence problem for the Euler and Navier-Stokes equations [8, 9, 10]. Slightly generalizing definitions of Tao [8, p. 220] and Cardona and Presas [11, pp. 3–4], we say that a class $\mathcal{C}$ of flows is $C^{k}$-universal with respect to another such class $\mathcal{D}$ if any flow in $\mathcal{D}$ admits an (equivariant) $C^{k}$ embedding into some flow in $\mathcal{C}$ (realizing the former as an invariant subset of the latter). If we fix
$\mathcal{C}=\{\text{linear flows on finite-dimensional Euclidean spaces}\},$
@@
our existence question can be restated equivalently as follows: what is the largest class of flows $\mathcal{D}$ with respect to which $\mathcal{C}$ is $C^{k}$-universal? We answer this question under the additional constraint that $\mathcal{D}\subset\mathcal{E}$, where $\mathcal{E}$ consists of flows on connected spaces that are either compact or contain at least one nonempty compact attractor.
1.1. Related work
Poincaré initiated the study of locally-defined, dimension-preserving linearizing embeddings (homeomorphisms and diffeomorphisms) near an equilibrium point [12, 13], leading to classical linearization theorems including those of Poincaré-Siegel [12, 14, 15], Sternberg [16, 17], Hartman-Grobman [18, 19] and Hartman [20]. Extensions to locally-defined, dimension-preserving, partial linearizations (in transverse directions) near general invariant manifolds were given in theorems including those of Pugh and Shub [21], Robinson [22], Takens [23], Sell [24, 25], and Sakamoto [26]. Such linearization results were extended to global ones valid on the basin of attraction of a stable hyperbolic equilibrium or limit cycle by Lan and Mezić [27] and, more generally, on the basin of attraction of a stable normally hyperbolic invariant manifold without boundary by Mezić [28] and with inflowing boundary by Eldering, Kvalheim, and Revzen [29].
In this paper we are interested in globally-defined, totally (not partially) linearizing $C^{k}$ embeddings that are possibly dimension-increasing. Linearizing $C^{k}$ embeddings are conceptually natural objects of study, since they are precisely the maps $C^{k}$-identifying nonlinear systems with invariant subsets of linear systems. However, previous literature has focused on related but different classes of maps.
To our knowledge, Carleman was the first to study linearizing embeddings without the dimension-preserving assumption [30, 31]. Both Carleman and recent control theory literature consider $C^{k\geq 1}$ embeddings $F\colon M\to\mathbb{R}^{n}$ satisfying the extra requirements that $M=\mathbb{R}^{m}$ and $F\colon\mathbb{R}^{m}\to\mathbb{R}^{n}$ has a linear left inverse $\mathbb{R}^{n}\to\mathbb{R}^{m}$, while only satisfying a weaker “linearizing” property than we require. Namely, the linear left inverse is merely required to send linear trajectories starting in the image of $F$ to nonlinear trajectories in $M$, with the image of $F$ not necessarily invariant under the linear flow. Belabbas, Chen, and Ko dubbed such embeddings $F$ “super-linearizations” and defined “strong super-linearizations” to be super-linearizations that are also linearizing in our sense [32, 33, 34, 35]. One necessary and sufficient condition for super-linearizability of a $C^{\infty}$ flow $\Phi$ on $\mathbb{R}^{m}$ was obtained by Claude, Fliess, and Isidori [36]. [^2] They showed that a $C^{\infty}$ flow $\Phi$ on $\mathbb{R}^{m}$ is super-linearizable in this sense if and only if $\text{span}\{x\mapsto\frac{d^{j}}{dt^{j}}\Phi^{t}(x)|_{t=0}\colon j\in\mathbb{N}\}$ is a finite-dimensional linear subspace of the real vector space of $C^{\infty}$ maps $\mathbb{R}^{m}\to\mathbb{R}^{m}$ [36] (see also Levine and Marino [37, Lem. 1]). Existence of a $C^{k}$ strong super-linearization for a flow $\Phi$ on $\mathbb{R}^{m}$ is also equivalent to its group of Koopman operators having a finite-dimensional invariant subspace of $C^{k}$ functions containing the state coordinate projections $(x_{1},\ldots,x_{m})\mapsto x_{i}$, a situation of interest in the recent “applied Koopman operator” literature [38]. Super-linearizations are also related to “polynomial flows” [39, p. 671].
On the other hand, Liu, Ozay, and Sontag [40, 41] recently obtained a necessary condition for existence of linearizing injective $C^{0}$ maps, called “one-to-one linear immersions” therein. These are equivalent to $C^{0}$ embeddings in situations considered in this paper (see Remark 4).
Previously, Mezić [6] introduced a necessary condition and a sufficient condition for existence of linearizing injective maps, called “faithful linear representations” therein. These conditions involve properties of the Koopman operator acting on spaces of $L^{2}(\mu)$ functions, where $\mu$ is a suitable measure, and the components of the linearizing injective maps are assumed to be in $L^{2}(\mu)$ but not necessarily continuous. Thus, the sufficient condition [6, Prop. 38] is not sufficient for linearizability by a $C^{0}$ embedding, but the necessary condition [6, Cor. 33] is also necessary for linearizability by a $C^{0}$ embedding if the state space is compact (or is replaced by a compact invariant set) and $\mu$ is a finite Borel measure.
@@ 1.2. Illustration of results
One obstruction to the existence of a linearizing $C^{0}$ embedding, if $M$ is connected, occurs if (1) has a compact attractor whose basin is not all of $M$ (Corollary 6). Hence we are led to study the existence of linearizing embeddings defined on attractor basins or other subsets, i.e., linearizing embeddings of $(X,\Phi)$ where $X\subset M$ is a subset invariant under $\Phi$ (by abuse of notation we still write $\Phi$ for $\Phi|_{\mathbb{R}\times X}$). Such a linearizing embedding $F\colon X\to\mathbb{R}^{n}$ satisfies (2) but is defined on $X$ rather than $M$.
Our results are divided among four cases: linearization by a $C^{k\geq 1}$ embedding when $X$ is an attractor basin or a compact invariant manifold, and linearization by a $C^{0}$ embedding when $X$ is an attractor basin or a compact invariant set.
@@
Our second main result (Theorem 2) concerns the case that $X$ is a compact invariant set (not assumed to be a manifold). It asserts that $(X,\Phi)$ is linearizable by a $C^{0}$ embedding if and only if $\Phi$ is a $1$-parameter subgroup of a $C^{0}$ torus action that is not too pathological (has finitely many “orbit types”). For example, all flows shown in Figure 1 are linearizable by $C^{0}$ embeddings.
Our third main result (Theorem 3) concerns the case that $X$ is the basin of attraction of an asymptotically stable compact invariant set $A\subset M$. [^3] Actually, $A$ need only be globally asymptotically stable within an invariant set $X\subset M$, as long as $X$ is “locally closed” in $M$ (see Remark 8). It asserts that $(X,\Phi)$ is linearizable by a $C^{0}$ embedding if and only if $(A,\Phi)$ is linearizable by a $C^{0}$ embedding (cf. Theorem 2) and $A$ has $C^{0}$ “asymptotic phase” for $\Phi$. Asymptotic phase as defined in §2.3 generalizes that associated with normally hyperbolic invariant manifolds [42, 43, 44, 29], which in turn generalizes that associated with hyperbolic periodic orbits [45, 46, 47]. Theorem 3 further asserts that linearizing $C^{0}$ embeddings (or, more generally, injective $C^{0}$ maps) are always proper and hence limit to $\infty$ near the basin boundary $\partial X$ (cf. [48, Prop. 1]), leading to the aforementioned statement (Corollary 6) that $(M,\Phi)$ is not linearizable by a $C^{0}$ embedding if $M$ is connected and there is a non-global compact attractor. Other corollaries of Theorem 3 are extensions of the Hartman-Grobman and Floquet normal form theorems, at the cost of allowing “extra” linearizing coordinates (Corollaries 4, 5).
Our fourth and final main result (Theorem 4) again concerns the case that $X$ is the basin of attraction of an asymptotically stable compact invariant set $A\subset M$, but instead concern linearizability by a $C^{k\geq 1}$ embedding. It asserts the following. If $(X,\Phi)$ is linearizable by a $C^{k}$ embedding, then $A\subset X$ must in fact be a $C^{k}$ embedded submanifold, $(A,\Phi)$ must be linearizable by a $C^{k}$ embedding (cf. Theorem 1), and $A$ must have $C^{k}$ asymptotic phase. Conversely, $(X,\Phi)$ is linearizable by a $C^{k}$ embedding if $A,X,\Phi$ have the preceding properties, $\Phi$ is the flow of a $C^{k}$ vector field (cf. Remark 10), and $\Phi$ is “transversely linearizable” near $A$ in a sense related to normal hyperbolicity (cf. Remarks 12, 13).
The remainder of the paper is organized as follows. Our general results and examples are in §2. Proofs are in §3. Closing remarks are in §4.
@@ 2. Results and examples
A flow on $X$ is a group action $\Phi\colon\mathbb{R}\times X\to X$ of $\mathbb{R}$ on $X$. If $\Phi$ is a flow on $X$, we say that a map $F\colon X\to\mathbb{R}^{n}$ is linearizing if there is $B\in\mathbb{R}^{n\times n}$ such that $F\circ\Phi^{t}=e^{Bt}\circ F$ for all $t\in\mathbb{R}$. We say that $\Phi$ is a 1-parameter subgroup of a group action $\Theta\colon H\times X\to X$ of a torus $H=T^{\ell}=\mathbb{R}^{\ell}/\mathbb{Z}^{\ell}$ on $X$ (a torus action) if there is $\boldsymbol{\omega}\in\mathbb{R}^{\ell}$ such that $\Phi^{t}=\Theta^{(\boldsymbol{\omega}t\mod 1)}$, where $\text{mod }1$” is applied entrywise.
@@ 2.1. The smooth compact case
With the aid of Theorem 1, the class of linearizable systems $(X,\Phi)$ as in its statement can be seen to be a subclass of the class of “non-Hamiltonian integrable systems” from the literature [49, Def. 2.6].
@@
Suppose $(X,\Phi)$ is either (i) a single equilibrium point, (ii) a single periodic orbit viewed as a constant-speed rotation on the circle $X=S^{1}$, or (iii) a single quasiperiodic torus viewed as a product of constant-speed rotations on the ($n$-)torus $X=T^{n}=S^{1}\times\cdots\times S^{1}$, a product of circles. It is clear (and well-known) that all of these cases are linearizable by $C^{\infty}$ embeddings. Let us compare this fact with Theorem 1. In all cases, $(X,\Phi)$ is a $1$-parameter subgroup of a $C^{\infty}$ torus action on $X$: the actions $\Theta$ can be respectively taken to be (i) the trivial action of $T^{1}$ on the equilibrium, (ii) the action of $S^{1}=T^{1}$ on itself, or (iii) the action of $T^{n}$ on itself.
@@
Let $X=S^{2}\subset\mathbb{C}\times\mathbb{R}\approx\mathbb{R}^{3}$ be the unit $2$-sphere centered at the origin. Let $\Theta$ be the $C^{\infty}$ action of the circle $S^{1}\approx\mathbb{R}/\mathbb{Z}$ on $\mathbb{C}\times\mathbb{R}$ defined by $\Theta^{h}(z,s)\coloneqq(e^{2\pi ih}z,s)$. This action leaves $S^{2}$ invariant because $\Theta$ simply rotates the sphere’s latitudinal circles. Thus, Theorem 1 implies that $(X,\Phi)$ with $\Phi^{t}\coloneqq\Theta^{t\mod 1}$ is linearizable by a $C^{\infty}$ embedding. This conclusion can also be seen directly since $\Phi$ is simply the restriction to $X\subset\mathbb{C}\times\mathbb{R}$ of the linear flow $(t,z,y)\mapsto(e^{2\pi it}z,y)$ on $\mathbb{C}\times\mathbb{R}\approx\mathbb{R}^{3}$. In anticipation of Proposition 1, note that $(X,\Phi)$ has two equilibria, each with Hopf index [50, pp. 133–134] equal to $1$.
@@
Proposition 1 and the Poincaré-Hopf theorem [51, p. 35] imply the following corollary, which is consistent with Examples 2, 3, 4, 5.
@@
The standard Poincaré-Hopf theorem and other results for $C^{\infty}$ manifolds also apply to $C^{1}$ manifolds since any $C^{1}$ manifold has a compatible $C^{\infty}$ structure [2, Thm 2.2.9].
@@
The following corollary follows directly from Corollary 2, Examples 2, 3, 4, 5, and the classification of surfaces [2, Thm 9.3.5].
@@
Theorem 1 gives a necessary and sufficient condition for linearizability by a $C^{k\geq 1}$ embedding. On the other hand, Proposition 1 and Corollaries 2, 3 give necessary conditions for the same, while the following proposition gives a sufficient condition. Proposition 2 is a “rigidity” statement about Koopman eigenfunctions, and its proof is independent of our other results. (Since $X$ is compact, assuming the existence of $F$ below is readily seen to be equivalent to assuming the existence of $n$ nowhere-zero $C^{k}$ Koopman eigenfunctions $X\to\mathbb{C}$ whose eigenvalues are rationally independent.)